OpenStudy (anonymous):

Describe a situation when you might not want to clear fractions in an equation.

OpenStudy (aravindg):

Ok take an example of expressing a line in slope intercept form if the equation is \[y-\dfrac{1}{2}x-3=0\] and if the question asks for finding slope we need not multiply the whole equation to remove fraction \(\dfrac{1}{2}\) .Instead just take 2nd and 3rd term to RHS \[y=\dfrac{1}{2}x+3\] slope=\(\dfrac{1}{2}\)

OpenStudy (aravindg):

@melissa98 does that help ?